OpenStudy (anonymous):

Given the functions f(x) = 7x + 13 and g(x) = x + 2, which of the following functions represents f[g(x)] correctly?

OpenStudy (anonymous):

f[g(x)] = 7x + 15 f[g(x)] = 7x + 27 f[g(x)] = 7x2 + 15 f[g(x)] = 7x2 + 27

OpenStudy (welshfella):